can you help me solve this problem?
The sum of their present ages is 40, The mother is 20 years older than the daughter. form an equation and solve it.

Answer by Cromlix(4381)

The mother is 20 years older than her daughter
Daughter = x
Mother = 20 + x
x + 20 + x = 40
2x + 20 = 40
2x = 40 - 20
2x = 20
x = 10
Daughter is 10 years old
Mother is 30 years old.
Hope this helps
:-)